July 23, 2020 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Time: 11:00 a.m. JLab time

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

  1. Short Term:
    1. Checks of monitoring comments of Batch 5-8
    2. Timing checks: 72123-72416 (2ns peaks), 72417-72435 (time offset) -> Mark
    3. BCAL gains calibration constants (nominal 8 batches grouped in 4) -> Karthik
    4. Pedestal checks (Primex) -> Igal
    5. Magnetic field in BCAL: check energy calibration -> Mark
    6. Richard Jones' work on G3vG4: resolution at 60 deg
    7. LED occupancy hole investigations
    8. BCAL Wiki reorganization -> All
  2. Long term items

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run Update
    1. FCAL
    2. BCAL
  4. Calibrations
  5. Monitoring
    1. BCAL LED
    2. BCAL (offline)
    3. FCAL (offline) statusIII
  6. Reconstruction
  7. Efficiency
  8. Simulations
    • Discussion (Jon): splitoffs in MC: it has been brought up across several meetings that hadronic splitoffs may not be modeled in MC well, thus caution is recommended when cutting on number of unused showers. Is there a quick study we could do to point towards this? For FCAL: exclusive omega sample may work. For BCAL: hyperons??
  9. Any other business


Attending: Tolga, Zisis, Sean, Mark, Elton, Colin, Susan, Varun, Jon, Churamani, Karthik, Igal, Rebecca, Matt

  1. Announcements
    1. Elton reported that the accelerator has issues (N. Linac) and beam will be delivered a week later than planned.
  2. Action Items
    1. Short Term
      1. Checks of monitoring comments of Batch 5-8 (see below)
      2. Mark made a log entry yesterday in connection to the timing checks. The average photon timing to 0 has been made for runs 71350 to 71591. He will check again after the next monitoring launch.
      3. Karthik is working on the "Group Batches" (=ones that have similar gains.) There are four of those. The last group (normal batches 7 and 8) are running and will be analyzed later today.
      4. Pedestal checks; use new gain calibrations, starting today.
      5. The magnetic field was off in the BCAL SciFi matrix for HGeant because is caused it to be slow; this does not matter for HDGeant4. Now the field is on for both. Mark verified that it does not affect the energy scale.
      6. There was a discussion connected to Richards' G3vsG4 work. Matt: are these the best observables? Acceptance is too complicated a variable. Shouldn't we use a monoenergetic particle gun and go back to basics? (Note HADR=1 and =4 are used only in HDGeant.) Sean: after event is reconstructed there is a big difference in the timing distributions between the two HADR options. It was agreed to do photon, pi+, pi-, p gun studies. WHO WILL DO THIS? Mark suggested that etapi-Delta++ is a good candidate to study the hadron timing. This differs from Alex's pi+pi-p (the rho).
      7. LEDs: Varun verified that SiPM M23 Layer 2 #6 is not functioning, by comparing that cell with old runs: the signal now is 1/2 of what it was when both SiPMs in the cell (#6 and #10) were firing. Elton: this is the first time we have a dead SiPM; we should consider knocking out this cell in ccdb. Electronics/FADC and amplifiers are functioning. It may be that the #6 cable is corroded or has a poor connection; this will be checked after the run. Also, Varun and Jon tool LED data towards the continued radiation damage study. Beni should be consulted as he is working on a similar study for the Start Counter.
      8. Elton mentioned that if more people contribute to the BCAL Wiki it will become a useful tool.
    2. Long term:
      1. Colin: Check this board; check the switch: maybe it is on the wrong scale. Check before the run.
      2. Mark: it is not obvious how to do this. Is it vital on the short term?
  3. Run Update
    1. FCAL: 10 bases need to be replaced; maybe tomorrow or next week.Colin and Mark are working on the hot checkout (HC). A new HV file will be created. They will run leD scripts and verify that the beam line is clear.
    2. BCAL: Need to sign off on the HC. A few more items left. Jon is checking the LED system.
  4. Calibrations: nothing to report.
  5. Monitoring
    1. BCAL LED: see Action Items above.
    2. BCAL (offline): nothing to report. Still checking over newest batches.
    3. FCAL (offline): red items show on slide 1; checked Matt's suggestions. Some runs were taken with LEDs at very different rates. To further pursue the study raw data needs to be staged to get ntuples-trees. To check the time offset compare the last 5 files of the run with the first 5.
  6. Reconstruction: nothing to report.
  7. Efficiency: Elton updated the group on his costheta efficiency work; this belongs to one of Sean's Task Forces. In the BCAL NIM plots look flat but now they do not look so nice. Assumption: low gamma is inefficient, high one is efficient. Comparisons with both photons in FCAL, both in BCAL, one in each calorimeter. MC vs data t-slope: FCAL slope 6, BCAL slope 2.5. Note that data and MC were not done with the same reconstruction version (MC has the wrong one). Elton looked at 2017 pristine run set. Mark: we need to model the mass and t-distribution better. Used HDG3 (HARD=1, 4) to see what the spread is. Elton: data 2017 to 2018 comparison is good; match to MC is not good. The central point of the study is that the theta distribution is related to the energy distribution in the lab. (opening angle has a 1-to-1 correspondence to invariant mass).
  8. Simulations: Jon reported that the Combo WG deals with this (hadronic splitoffs). Mark did a study with omega and will repeat it in two months. Sean: effect is final-state dependent. We should add this to our list of priorities: objective: determine # of showers and # of splitoffs? Or, model Unused Energy (may be easier that no of showers). Sean: are the splitoffs related at all to the presence of the magnetic field?
  9. Any other business: Sean: covariance matrices for the photons. Mark did them 4 years ago. They should be redone NOW before reconstruction: try G3 (HADR1,4) and G4.